OpenStudy (anonymous):

Here's a question for AriPotta, and for all of you! It may seem hard, but it's not, and I actually think it's quite fun, if you need help, just ask :) The total of three cousin's ages is 48. Bobby is half as old as Jim and 4 years older than Bob. How old are the cousins?

OpenStudy (anonymous):

One is 11, one is 22 and one is 15

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

Anyone got it?

OpenStudy (anonymous):

13, 26, 9

OpenStudy (anonymous):

good, dont tell how you got it tho...I want it to be a surprise

OpenStudy (anonymous):

\(a=\frac12b=c+4\) which means we can write their ages all in terms of a: \(b=2a\), \(c=a-4\). The sum of their ages \(a+b+c\) is \(48\) i.e. \(a+2a+a-4=48\); solve for \(a\) by combining like terms to yield \(4a-4=48\) which reduces to \(a-1=12\), so \(a=13\). Solve for our other ages \(b\) and \(c\) straightforwardly to yield \(b=2(13)=26\) and \(c=13-4=9\)

OpenStudy (anonymous):

gosh...lol thats a lot of numbers there lol

OpenStudy (anonymous):

Bobby is 13, Jim is 26, and Bob is 9.

OpenStudy (anonymous):

good